I need to solve this problem and show solution used....HELP!!!!
The length of a rectangle is 2 in more than twice its width . If the perimeter of the rectangle is 34 in find the dimensions of the rectangle.
thanks alot
Heather

Answer by checkley71(8403)

L=2W+2
PERIMETER=2L+2W
34=2(2W+2)+2W
34=4W+4+2W
6W=34-4
6W=30
W=30/6
W=5 ANSWER.
L=2*5+2
L=10+2
L=12 ANSWER.
PROOF
2*12+2*5=34
24+10=34
34=34